Joe the grocer bought some oranges at the cost of 3 for 48 cents, then twice as many at the cost of 4 for 58 cents. In order to make a profit of 25%, he must sell them all at a price of 4 for m cents. The value of m in cents is:
A) 60
B) 75
C) 80
D) 62
E) 63

Answer by richwmiller(17219)

6 at 3 for 48 2*48=96
12 at 4 for 58 3*58 =174
270/18=15 each
4 *15=60
A) 60